# Multi-task neural diffusion processes for uncertainty-quantified wind power prediction

### v1

@@ Abstract
Uncertainty-aware wind power prediction is essential for grid integration and reliable wind farm operation. We apply neural diffusion processes (NDPs)—a recent class of models that learn distributions over functions—and extend them to a multi-task NDP (MT-NDP) framework for wind power prediction. We provide the first empirical evaluation of NDPs in real supervisory control and data acquisition (SCADA) data. We introduce a task encoder within MT-NDPs to capture cross-turbine correlations and enable few-shot adaptation to unseen turbines. The proposed MT-NDP framework outperforms single-task NDPs and GPs in terms of point accuracy and calibration, particularly for wind turbines whose behaviour deviates from the fleet average. In general, NDP-based models deliver calibrated and scalable predictions suitable for operational deployment, offering sharper, yet trustworthy, predictive intervals that can support dispatch and maintenance decisions in modern wind farms.
@@
Few-shot inference , Probabilistic predictions , Renewable energy , SCADA data
journal: Energy and AI
[…]
[label1]organization=Department of Statistical Science, University College London, addressline=Gower Street, city=London, postcode=WC1E 6BT, country=United Kingdom
[…]
[label2]organization=Department of Statistics, Athens University of Economics and Business, addressline=28is Oktovriou 76, city=Athens, postcode=104 34, country=Greece
@@ 1 Introduction
Wind energy has become a cornerstone of the global transition to clean power. As wind power capacity expands worldwide, ensuring reliability and minimising downtime are critical to both energy security and the financial viability of wind farms. Uncertainty-aware and accurate prediction of wind power output is vital, as it supports grid stability, market participation, and proactive maintenance planning. Beyond energy balancing, uncertainty-aware forecasting also reduces operational uncertainty for wind farm operators, enabling more efficient maintenance scheduling and reducing costly unplanned downtime. This is especially important given that operation and maintenance costs represent a significant share of total expenditure, with unexpected failures making up the largest component [1, 2].
Supervisory control and data acquisition (SCADA) systems provide a low-cost and widely available source of wind turbine data. They capture environmental and operational variables with high frequency, making them invaluable for prediction applications. However, their use is complicated by measurement noise, turbine downtime, and limited public availability [3, 4]. A practical prediction model for SCADA data must therefore satisfy several requirements: (i) scalability to large SCADA datasets covering multiple years of wind farm operation, (ii) the ability to exploit the high-dimensional feature space recorded in SCADA systems, (iii) probabilistic power forecasts with calibrated predictive densities rather than only point estimates, and (iv) the ability to generalise across turbines with limited historical records.
Several modelling approaches have been proposed to meet these requirements. Quantile regression methods generate predictive intervals directly and have been applied in spatio-temporal and deep learning settings for wind power forecasting [5, 6, 7]. Mixture density networks have been proposed to represent multimodal output distributions [8], while generative moment matching networks provide distribution-free probabilistic scenarios without assuming specific functional forms [9]. These methods enhance flexibility in modelling uncertainty, but generally lack mechanisms for transfer learning across wind turbines and can be computationally demanding when scaled to large SCADA datasets.
Gaussian processes (GPs) have been widely applied in wind power forecasting due to their ability to provide calibrated predictive distributions [10]. For example, heteroscedastic GP models have been used to capture input-dependent variance [11]. Other studies have incorporated turbine operational variables to improve forecast accuracy [12]. GPs have also been combined with numerical weather prediction outputs to enhance predictive performance [13]. Despite these strengths, the cubic scaling of kernel computations ($O(n^{3})$ in the number of data points, $n$) makes GPs computationally expensive on large SCADA datasets, and their strict Gaussian assumptions often limit performance in highly non-linear regimes [14].
Neural network (NN)-based approaches, such as probabilistic multi-layer perceptrons (MLPs), provide scalability and enable transfer learning across turbines, but they still model predictive uncertainty by explicitly predicting the mean and variance under a Gaussianity assumption [15]. Neural processes (NPs) combine the predictive power of NNs with the uncertainty estimation of GPs using an encoder–decoder architecture [16]. However, they often suffer from underfitting and limited consistency between context and target distributions [17]. Extensions such as multi-task neural processes demonstrate the potential to capture correlations across tasks [18], a particularly relevant issue for wind power prediction, since wind turbines on the same wind farm are typically exposed to similar environmental conditions.
Recent advances in generative modelling have introduced diffusion models, which iteratively corrupt and denoise data to learn complex probability distributions [19, 20]. These are state-of-the-art models in the domain of image generation [21]. Neural diffusion processes (NDPs) extend this framework to distributions over functions, avoiding Gaussianity assumptions while retaining scalability through the neural architecture. NDPs have been shown to emulate GPs well on synthetic data and outperform NPs [22]; however, their applications to real-world datasets remain unexplored in the regression setting.
We are the first to apply NDPs to SCADA data and to develop the methodology that allows adaptation to the multi-task setting. Specifically, we develop a novel multi-task NDP (MT-NDP) framework that incorporates a task encoder to capture correlations across turbines, enabling knowledge transfer and improved predictive performance when data for individual wind turbines are limited. The key advances of this paper are: (i) the first empirical evaluation of NDPs on operational SCADA data, showing accuracy exceeding that of GPs while delivering better-calibrated uncertainty estimates and scalability to higher-dimensional inputs; (ii) the extension of NDPs to a multi-task setting via a task encoder, which leverages cross-turbine correlations and consistently outperforms single-task variants; (iii) demonstration of the ability of MT-NDPs to perform few-shot generalisation, adapting to unseen turbines with only a small number of context points; and (iv) empirical evidence of scalability and robustness, with predictions that remain uncertainty-aware and suitable for decision-making in real-world wind farm operations.
The remainder of the paper is organised as follows. Section 2 reviews the benchmark models and introduces the background for NDPs. Section 3 describes the SCADA wind farm data. Section 4 presents the NDP framework and our multi-task extension (MT-NDP). Section 5 outlines the experimental design, and Section 6 reports the main findings. Finally, Sections 7 and 8 provide discussion and conclusions.
2 Preliminaries
In this section, we present the foundational models that support our framework or serve as benchmarks for evaluating the proposed multi-task extension. Section 2.1 covers GPs, Section 2.2 reviews diffusion probabilistic models, and Section 2.3 describes NDPs and provides the theoretical basis for the multi-task extension introduced in Section 4.
2.1 Gaussian Processes (GPs)
A GP $f:\mathbb{R}^{D}\rightarrow\mathbb{R}$ is a stochastic process such that, for any finite collection of inputs $X=\{x_{i}\}_{i=1}^{N}\subset\mathbb{R}^{D}$, the vector of function evaluations $f(X)=[f(x_{1}),\ldots,f(x_{N})]^{T}$ follows a multivariate normal distribution [10]. GPs satisfy the Kolmogorov extension theorem, which guarantees that all finite-dimensional marginals are mutually consistent under permutation and marginalisation [23]. This provides a principled Bayesian framework for modelling distributions over functions, enabling exact inference with predictive uncertainty. In our experiments presented in Section 6, we use GP regression as a benchmark model, so we briefly outline its formulation here.
$f(X)\sim\mathcal{N}(\mu(X),K(X,X)),$ (1)
$E(y^{*}|y)$ $=\mu(X^{*})+K(X^{*},X)A^{-1}(y-\mu(X)),$ (2)
$V(y^{*}|y)$ $=K(X^{*},X^{*})-K(X^{*},X)A^{-1}K(X,X^{*}).$ (3)
For our experiments, we use the radial basis function (RBF) kernel:
2.2 Diffusion probabilistic models (DPMs)
DPMs were introduced as a class of generative models inspired by non-equilibrium thermodynamics [19]. The central idea is to define a forward process, in which the structure of the data distribution is gradually destroyed by the iterative addition of Gaussian noise, and a corresponding reverse process that learns to invert this corruption.
The de-noising diffusion probabilistic model (DDPM) [20] simplified and stabilised training and has become the standard formulation. Following [20, 22], the forward process begins from the input data distribution $q(s_{0})$, where $s_{0}\in\mathbb{R}^{D}$ denotes a $D$-dimensional data vector. It defines a fixed Markov chain over the sequence $s_{0:T}=\{s_{0},\ldots,s_{T}\}$ with density
$q(s_{0:T})=q(s_{0})\prod_{t=1}^{T}q(s_{t}\mid s_{t-1})$ (5)
and conditional distributions
where $\{\beta_{t}\in(0,1)\}_{t=1}^{T}$ is a variance schedule controlling the magnitude of injected noise and $I_{D}$ is the $D\times D$ identity matrix. After $T$ steps, the distribution approaches Gaussian noise, $q(s_{T})\approx\mathcal{N}(0,I)$. Early implementations used a linear variance schedule, while later work demonstrated that a cosine schedule improves performance by distributing noise more evenly across diffusion steps [24]. A variance schedule defines the cumulative noise parameter
$\bar{\alpha}_{t}=\prod_{j=1}^{t}(1-\beta_{j}),$ (7)
The reverse process is intractable in closed form, but can be approximated by an NN that predicts the injected noise. The mean of the reverse transition is parameterised as
where $\alpha_{t}=1-\beta_{t}$, $\bar{\alpha}_{t}=\prod_{j=1}^{t}\alpha_{j}$, and $\epsilon_{\theta}:\mathbb{R}^{D}\times\{1,\dots,T\}\to\mathbb{R}^{D}$ is a neural network that predicts the injected noise vector. Training is carried out by minimising the score-matching loss
This framework enabled diffusion models to achieve state-of-the-art generative performance, rivalling generative adversarial networks (GANs) and autoregressive models. Building on DDPMs, Repaint [21] adapted the approach to image inpainting by conditioning on context points during the reverse process, without altering the pre-trained DDPM. This method achieved high-quality and diverse reconstructions under challenging conditions, outperforming autoregressive and GAN-based baselines.
In parallel, NPs were proposed as a stochastic process model combining neural networks with an encoder–decoder architecture [16]. NPs learn to map a set of context points into a latent task representation, which is then used by a decoder to generate predictions at target inputs. This provides scalability and the ability to adapt to new tasks with only a few observations, but standard NPs often underfit and produce poorly calibrated uncertainty [17]. Latent and attentive variants improve correlations between targets, but still fall short of fully consistent Bayesian behaviour [22].
Neural diffusion processes (NDPs) unify these two developments, drawing on diffusion-based generative modelling to overcome the Gaussian assumptions of NPs while retaining their scalability and adaptability. NDPs extend diffusion models from distributions over data to distributions over functions [22], thereby enabling uncertainty-aware regression.
2.3 Neural diffusion processes (NDPs)
@@
The core idea of NDPs is to replace the fixed data vector $s_{0}$ in DDPMs with function evaluations $(X,y)$. Here $X\in\mathbb{R}^{N\times D}$ is the matrix of input locations and $y\in\mathbb{R}^{N}$ the corresponding outputs. In the forward process, Gaussian noise is gradually added to the function values $y$, while the inputs $X$ remain fixed. After $T$ steps, the corrupted outputs resemble white noise. The reverse process is parameterised by an NN that predicts the noise injected at each step, thereby enabling recovery of the original function values. Training minimises a denoising score-matching objective, analogous to that of DDPMs, as described in Section 4.1.3.
Compared to GPs, NDPs avoid restrictive Gaussianity assumptions by learning arbitrary distributions over functions. Compared to NPs, they offer better-calibrated predictive distributions and a more consistent treatment of context and target points. On synthetic benchmarks, NDPs have been shown to emulate GP behaviour while scaling more effectively through neural architectures. In Section 4, we provide a detailed description of NDPs, together with illustrations of the forward and reverse processes, and introduce our proposed multi-task extension (MT-NDP).
3 Dataset description
We use SCADA data from the six Senvion MM92 wind turbines at Kelmarsh wind farm in the UK [25]. The dataset spans January 2016 to July 2021 at a ten-minute resolution, comprising more than 1.7 million records across 110 variables, including wind speed, wind direction, power output, and component temperatures. Each variable is reported as mean, minimum, maximum, and standard deviation within the ten-minute bins.
An operational status and events file was used to filter the SCADA data and ensure consistent modelling of normal turbine behaviour. These logs provide detailed information on turbine operating states, including technical failures as well as operational or environmental standbys and warnings. To enable accurate modelling, all data associated with out-of-control conditions were removed so that model training was based solely on stable operating periods. Specifically, records corresponding to standbys, warnings, and operational stops were excluded. In addition, data from the week preceding each forced outage was discarded to reduce the likelihood of capturing deteriorating states. The resulting data elimination process is also described and illustrated in [15].
Figure 1: SCADA data for the six turbines from the Kelmarsh wind farm [25] after removing standbys and warnings using the operational status and events file. (a) Wind power curves showing mean power output (kW) by mean wind speed (m/s). (b) Joint distribution of mean wind speed (m/s) and mean wind direction (°) for turbine 1 (left), and mean power output (kW) by mean wind speed (m/s) and wind direction (°) for turbine 1 (right).
Figure 2: SCADA data for the six turbines from the Kelmarsh wind farm [25] after removing standbys and warnings using the operational status and events file. (a) Mean power output (kW) by mean nacelle temperature (°C). (b) Mean power output (kW) by mean transformer temperature (°C).
4 Methodology
We present the methodology underlying our proposed multi-task NDP framework. We first formalise the NDP in the regression setting, where the goal is to model a distribution over functions $f:\mathbb{R}^{D}\to\mathbb{R}$. Each training instance corresponds to a sampled function evaluated at a finite set of input locations $X\in\mathbb{R}^{N\times D}$ with corresponding outputs $y=f(X)\in\mathbb{R}^{N}$. By training across many such functions, the model learns an empirical covariance structure that captures correlations between inputs and outputs, enabling calibrated predictive distributions even in regions with sparse data.
We outline the forward and reverse processes, and the sampling procedure, highlighting parallels with the DDPM framework. We then introduce our multi-task extension, which incorporates a task encoder to capture cross-turbine dependencies in wind farms.
4.1 Neural diffusion processes (NDPs)
@@
We begin by formalising the training data used by NDPs. Each training instance corresponds to a sampled function $f_{i}:\mathbb{R}^{D}\to\mathbb{R}$ evaluated at a finite set of inputs $x_{i}\in\mathbb{R}^{N\times D}$ with outputs $y_{i}=f_{i}(x_{i})\in\mathbb{R}^{N}$. The overall training dataset is therefore
@@
4.1.1 Forward process
Let $x_{0}=x$ and $y_{0}=y$. Then the NDPs gradually add noise following
@@ 4.1.1 Forward process
for each timestep $t$ where $\beta_{t}$ is the noise from the variance schedule [22]. Please note the similarities to Eq. 6, where now, rather than just modelling the input data $s_{t}$, we explicitly differentiate between the input locations $x_{t}$ and outputs $y_{t}$. This corresponds to adding Gaussian noise to the function values $y_{t}$ while keeping the input locations $x_{t}$ fixed for all $t$. By the final timestep ($t=T$), the function values $y_{t}$ will look like Gaussian noise from $\mathcal{N}(0,1)$. Fig. 3 illustrates the forward diffusion process, where Gaussian noise is gradually added to the outputs using a cosine variance schedule [24]. As $t$ increases, the data loses structure and converges towards white noise.
Figure 3: Illustration of the forward diffusion process. Gaussian noise is added incrementally at each timestep under a cosine variance schedule, progressively degrading the structure of the data until it converges to white noise. GP= Gaussian process, CI= Confidence interval
4.1.2 Reverse Process
NDPs use a neural network that learns to de-noise the corrupted function values, in a very similar manner to the DDPMs of [20]. The input locations $x_{t}$ are not corrupted, while the $y_{t}$ are. In fact, the model will use the uncorrupted $x_{t}$ to help improve the de-noising model. The parametrised kernel is of the form [22]:
@@ 4.1.2 Reverse Process
where the noise model $\epsilon_{\theta}:\mathbb{R}^{N\times D}\times\mathbb{R}^{N}\times\mathbb{R}\;\to\;\mathbb{R}^{N}$ takes as input the non-corrupted $x_{t}$ as well as the corrupted $y_{t}$, and the timestep $t$. As previously, $\alpha_{t}=1-\beta_{t}$ and $\bar{\alpha}_{t}=\prod_{j=1}^{t}\alpha_{j}$. Eq. 14 differs from the previous DDPM formulation Eq. 9 since it uses just the $s_{t}$ as input to the noise model rather than explicitly using uncorrupted $x_{t}$ and corrupted $y_{t}$. The goal of this noise model is to predict the noise added to $y_{0}$ to get $y_{t}$ in line with the DDPM model in [20].
4.1.3 Objective
The objective is very similar to Eq. 10 but with $s_{t}$ replaced by $x_{t}$ and $y_{t}$:
@@ 4.1.3 Objective
4.1.4 Sampling
Unconditional sampling: Given some input locations $x$, we can sample unconditionally from the trained model prior by starting with random noise and following the reverse process for $T$ timesteps. Specifically, we initialise $y_{T}\sim\mathcal{N}(0,I)$ and then apply the parametrised kernel for $t=T,\dots,1$. At each step, the neural network predicts the noise added in the forward process, and new noise is reintroduced. This procedure prevents the trajectories from collapsing to a local maximum and ensures coverage of the full distribution.
Unconditional samples already yield realistic outputs that resemble the true data distribution. In practice, however, if the training dataset is highly varied, unconditional sampling may not capture sufficient structure. For wind turbine data, where the relationships between inputs and power output are similar across turbines, the prior alone often produces plausible results.
Conditional sampling: One of the main strengths of NDPs, as with NPs, is their ability to quickly adapt with only a small amount of data. Conditional sampling is achieved by reinjecting context points at each timestep, using a slight adaptation of the Repaint algorithm [21]. This guides the reverse process to remain consistent with the context set.
@@ 4.1.4 Sampling
Repeating this for $t=T,T-1,\dots,1$ means that, at each backward step, the context points steer the process in the desired direction. Initially, when $y^{*}_{t}$ is close to pure noise, the context points already contain useful structure. This pulls the reverse trajectories towards functions consistent with the context set, leading to samples that better fit the true data.
In Fig. 4, we compare unconditional and conditional sampling. Even at early timesteps, the conditional model is visibly guided by the context points, producing a closer match to the true data throughout the process.
Figure 4: Comparison of unconditional and conditional reverse diffusion sampling. We sampled 10 reverse trajectories per input location, plotting the mean of these. Conditional sampling provides a closer fit to the true data across all timesteps. GP= Gaussian process
To illustrate the dynamics more clearly, Fig. 5 depicts the forward and reverse trajectories for a single input point. In the forward process (blue), Gaussian noise gradually corrupts the true value. In the reverse process (red), starting from pure noise, the trajectories converge back to the true value, while uncertainty narrows as the timestep decreases.

### v2

@@ Abstract
Neural diffusion processes provide a scalable, non-Gaussian approach to modelling distributions over functions, but existing formulations are limited to single-task inference and do not capture dependencies across related tasks. In many multi-task regression settings, jointly modelling correlated functions and enabling task-aware conditioning is crucial for improving predictive performance and uncertainty calibration, particularly in low-data regimes. We propose multi-task neural diffusion processes, an extension that incorporates a task encoder to enable task-conditioned probabilistic regression and few-shot adaptation across related functions. The task encoder extracts a low-dimensional representation from context observations and conditions the diffusion model on this representation, allowing information sharing across tasks while preserving input-size agnosticity and the equivariance properties of neural diffusion processes. The resulting framework retains the expressiveness and scalability of neural diffusion processes while enabling efficient transfer to unseen tasks. Empirical results demonstrate improved point prediction accuracy and better-calibrated predictive uncertainty compared to single-task neural diffusion processes and Gaussian process baselines. We validate the approach on real wind farm data appropriate for wind power prediction. In this high-impact application, reliable uncertainty quantification directly supports operational decision-making in wind farm management, illustrating effective few-shot adaptation in a challenging real-world multi-task regression setting.
@@
Few-shot inference, uncertainty quantification, probabilistic regression, renewable energy, wind power prediction
@@ 1 Introduction
Diffusion probabilistic models have emerged as a powerful approach for learning complex distributions via iterative denoising [sohldickstein2015deepunsupervisedlearningusing, ho2020denoisingdiffusionprobabilisticmodels]. Neural diffusion processes (NDPs) are a class of probabilistic models that learn distributions over functions by extending diffusion-based generative modelling to function space [dutordoir2023neural]. They operate by iteratively corrupting and denoising function values, avoiding explicit Gaussian assumptions while retaining scalability through neural network (NN) parameterisations. In the regression setting, NDPs support uncertainty-aware prediction and allow conditioning on context observations at inference time via conditional inference [dutordoir2023neural]. These properties make NDPs a flexible foundation for probabilistic regression in complex and non-linear settings.
We introduce multi-task neural diffusion processes (MT-NDPs), an extension of NDPs for regression problems involving collections of related tasks. Many real-world regression settings are inherently multi-task, where functions share common structure while exhibiting task-specific variation, such as multi-sensor measurements [alvarez2012kernels], spatio-temporal environmental data [wang2016locally], and wind turbines operating within a wind farm [Fiocchi_2025, Hadjoudj2023AOpportunities, wes-8-893-2023]. Jointly modelling such tasks improves data efficiency and predictive performance, particularly in low-data regimes [caruana1997multitask]. However, existing NDPs are restricted to single-task inference and cannot exploit cross-task correlations, motivating a principled multi-task extension.
Prior work on multi-task probabilistic regression has explored several approaches for sharing information across different tasks. Multi-task Gaussian processes (GPs) provide a Bayesian framework for joint inference via structured cross-task covariance models, including autoregressive formulations that handle incomplete observations across tasks [bonilla2007multi, alvarez2012kernels, requeima2019gaussian]. While effective, these methods scale poorly to large datasets and depend strongly on kernel and covariance design [Rasmussen2006GaussianLearning]. Neural alternatives such as neural processes (NPs) improve scalability by learning task representations from context sets, and multi-task extensions enable transfer learning across related functions [garnelo2018neuralprocesses, kim2019attentiveneuralprocesses, kim2022multitaskneuralprocesses]. However, these models often suffer from underfitting and inconsistent uncertainty calibration. NDPs address these limitations by learning expressive, non-Gaussian distributions over functions [dutordoir2023neural], but existing formulations are restricted to single-task inference.
Wind energy systems provide a representative and practically relevant example of multi-task regression, where multiple wind turbines operate under similar environmental conditions while exhibiting turbine-specific behaviour. Accurate and uncertainty-aware wind power prediction is a high-impact application and it is essential for grid stability and maintenance planning, particularly given that operation and maintenance costs account for a substantial fraction of total wind farm expenditure [en13123132, Hadjoudj2023AOpportunities]. Prior work has explored a range of probabilistic and machine learning approaches for wind power prediction, including GPs [2026probabilistic, rogers2020, pandit2020, chen2014], quantile regression [yu2019probabilistic, yu2021, zhou2022], mixture density networks [zhang2020], and NN–based probabilistic models [Fiocchi_2025]. While these methods provide varying degrees of flexibility and uncertainty quantification, many either scale poorly to large datasets or lack principled mechanisms for sharing information across turbines.
To address these challenges, we propose MT-NDPs, which incorporate a task encoder that maps context observations to a low-dimensional task representation and conditions the diffusion denoiser on this representation. This design enables information sharing across related tasks while preserving the set-structured, scalable, and uncertainty-aware properties of NDPs. As a result, MT-NDPs support few-shot adaptation to unseen tasks by conditioning on limited task-specific context at inference time.
We evaluate the proposed framework on wind power prediction from supervisory control and data acquisition (SCADA) data, assessing predictive accuracy, uncertainty calibration, and generalisation to unseen wind turbines under varying context sizes. The results show that MT-NDPs consistently outperform single-task NDPs and GP-based baselines, particularly in regimes with limited task-specific data.
The remainder of the paper is organised as follows. Section 2 introduces NDPs and their core properties. Section 3 presents the proposed multi-task extension. Section 4 describes the dataset and experimental setup, and Section 5 reports the empirical results. Section 6 discusses implications and limitations, and Section 7 concludes.
2 Background: Neural diffusion processes (NDPs)
This section aims to provide an overview of the key concepts used throughout the manuscript. Diffusion probabilistic models learn complex distributions by gradually corrupting data with noise and training a NN to reverse this process [ho2020denoisingdiffusionprobabilisticmodels]. NDPs adapt this framework from finite-dimensional data vectors to function values, enabling probabilistic regression with uncertainty quantification [dutordoir2023neural]. For completeness, we summarise the diffusion framework in Appendix B. By operating directly in function space, NDPs avoid the restrictive Gaussian assumptions of GPs while retaining scalability through neural architectures. Compared to NPs, NDPs yield better-calibrated predictive distributions and a more consistent treatment of context and target points. On synthetic benchmarks, NDPs have been shown to closely emulate GP behaviour while scaling more effectively to larger datasets [dutordoir2023neural].
@@ 2 Background: Neural diffusion processes (NDPs)
The core idea of NDPs is to replace the fixed finite-dimensional data vector used in standard diffusion models with function evaluations $(X,y)$. Here $X\in\mathbb{R}^{N\times D}$ is the matrix of input locations and $y\in\mathbb{R}^{N}$ the corresponding outputs. In the forward process, Gaussian noise is gradually added to the function values $y$, while the inputs $X$ remain fixed. After $T$ steps, the corrupted outputs resemble white noise. The reverse process is parameterised by a NN that predicts the noise injected at each step, thereby enabling recovery of the original function values.
@@
We now formalise this intuition by introducing the training data representation used by NDPs. Each training instance corresponds to a sampled function $f_{i}:\mathbb{R}^{D}\to\mathbb{R}$ evaluated at a finite set of inputs $x_{i}\in\mathbb{R}^{N\times D}$ with outputs $y_{i}=f_{i}(x_{i})\in\mathbb{R}^{N}$. The overall training dataset is therefore
@@
Forward process. Let $x_{0}=x$ and $y_{0}=y$. Then the NDPs gradually add noise following
@@
for each timestep $t$ where $\beta_{t}$ is the noise from the variance schedule [dutordoir2023neural]. This forward process is directly analogous to the standard diffusion formulation (Eq. 20 in Appendix B), but adapted to function-space regression by explicitly separating input locations $x_{t}$ from function values $y_{t}$. This corresponds to adding Gaussian noise to the function values $y_{t}$ while keeping the input locations $x_{t}$ fixed for all $t$. By the final timestep ($t=T$), the function values $y_{t}$ will look like Gaussian noise from $\mathcal{N}(0,1)$.
Reverse process. The reverse process is parameterised by a NN that predicts the noise injected during the forward diffusion process [ho2020denoisingdiffusionprobabilisticmodels]. The reverse transition is given by [dutordoir2023neural]:
@@
where the noise model $\epsilon_{\theta}:\mathbb{R}^{N\times D}\times\mathbb{R}^{N}\times\mathbb{R}\;\to\;\mathbb{R}^{N}$ takes as input the non-corrupted $x_{t}$ as well as the corrupted $y_{t}$, and the timestep $t$. As in standard diffusion models, $\alpha_{t}=1-\beta_{t}$ and $\bar{\alpha}_{t}=\prod_{j=1}^{t}\alpha_{j}$.
Training objective. Training proceeds by minimising a denoising score-matching objective,
@@
Unconditional sampling. Given some input locations $x$, we can sample unconditionally from the trained model prior by starting with random noise and following the reverse process for $T$ timesteps. Specifically, we initialise $y_{T}\sim\mathcal{N}(0,I)$ and then apply the parametrised kernel for $t=T,\dots,1$. At each step, the NN predicts the noise added in the forward process, and new noise is reintroduced.
Conditional sampling. One of the main strengths of NDPs, as with NPs, is their ability to quickly adapt with only a small amount of data. Conditional sampling is achieved by reinjecting context points at each timestep, using a slight adaptation of the Repaint algorithm [lugmayr2022repaintinpaintingusingdenoising]. This guides the reverse process to remain consistent with the context set.
@@
Repeating this for $t=T,T-1,\dots,1$ means that, at each backward step, the context points steer the process in the desired direction.